# SPF-Portrait: Towards Pure Portrait Customization with Semantic Pollution-Free Fine-tuning

### v1

@@ Abstract
While fine-tuning pre-trained Text-to-Image (T2I) models on portrait datasets enables attribute customization, existing methods suffer from Semantic Pollution that compromises the original model’s behavior and prevents incremental learning. To address this, we propose SPF-Portrait, a pioneering work to purely understand customized semantics while eliminating semantic pollution in text-driven portrait customization. In our SPF-Portrait, we propose a dual-path pipeline that introduces the original model as a reference for the conventional fine-tuning path. Through contrastive learning, we ensure adaptation to target attributes and purposefully align other unrelated attributes with the original portrait. We introduce a novel Semantic-Aware Fine Control Map, which represents the precise response regions of the target semantics, to spatially guide the alignment process between the contrastive paths. This alignment process not only effectively preserves the performance of the original model but also avoids over-alignment. Furthermore, we propose a novel response enhancement mechanism to reinforce the performance of target attributes, while mitigating representation discrepancy inherent in direct cross-modal supervision. Extensive experiments demonstrate that SPF-Portrait achieves state-of-the-art performance.
{NoHyper}* Co-first authors. {NoHyper} Work done during internship at KwaiVGI, Kuaishou Technology.{NoHyper}🖂 Corresponding authors.
1 Introduction
Fine-tuning a pre-trained Text-to-Image (T2I) model [43, 42, 45] has been widely recognized as an efficient and resource-saving solution for text-driven portrait customization [21, 15], especially for its extensive applications and substantial commercial value [31, 49, 25, 54, 29]. It adapts the model to understand personalized target attributes on an expansive text-image portrait dataset. An ideal fine-tuning should approach incremental learning [23], thereby maintaining the model’s original capabilities and continuous scalability [38, 36], while avoiding catastrophic forgetting [53, 11, 27]. This necessitates purely capturing the target semantics during fine-tuning, which manifests in the results as introducing portrait differences only through target attributes while maintaining consistency in unrelated attributes with the original portrait. However, as shown in Fig. LABEL:fig:highlight, although prevalent naive fine-tuning can achieve target attributes, it significantly changes the portrait identity, posture, background, and other attributes, disrupting the original model’s behavior. This is because the target semantics entangling with redundant features from the additional dataset during fine-tuning, resulting in an impure target response [37]. We refer to this phenomenon as “Semantic Pollution”, which is detrimental and often ignored. To address this issue, we propose SPF-Portrait, the first method to our knowledge that customizes attributes while eliminating semantic pollution in text-driven portrait customization.
@@ 1 Introduction
One line of previous research related to our work is PEFT-based methods [19, 10, 55, 3, 34]. For example, LoRA [19] and its variants [10, 55, 3] encourage the model to focus on the more relevant aspects of the target semantics by reducing trainable parameters, thus minimizing disruption. Orthogonal Fine-Tuning methods [40, 34] improve the LoRA on the model stability, further reducing disruption. However, their exclusive reliance on diffusion loss to implicitly model the joint distribution, rather than understanding disentangling semantics, making it challenging to preserve the original behavior. Another line of work [6, 58, 7, 50, 5, 35, 33, 22] aims to purify the understanding of text embeddings and decouple attributes from each other. For instance, Magnet [58] and TEBopt [7] focus on decoupling at the embedding level, while Tokencompose [50] apply regularization techniques to the attention mechanisms in model architecture for text understanding. However, these studies primarily improve the independence of target semantics in instance-level generation (e.g., a cat, a dog), but pay less attention to refined attributes, such as skin texture, freckles, and hairstyles of portraits.
@@
In this paper, we begin by defining two key objectives regarding the achievement of desired results in text-driven portrait customization: (1) Keeping portrait attributes unrelated to the target semantics consistent with the original model’s behavior; (2) Ensuring an effective response to the target attributes. To this end, we propose SPF-Portrait, a dual-path contrastive learning pipeline. We introduce the frozen original model to serve as the standard of original behavior for conventional fine-tuning path. To align with the original performance, we extract and constrain attention features and UNet features from the corresponding cross-attention layers in contrastive paths. We design a novel Semantic-Aware Fine Control Map (SFCM) that accurately identifies target response regions to spatially guide the alignment of these intermediate features. This alignment process precisely aligns irrelevant attributes, avoiding suppression of target response or over-alignment. Moreover, we introduce a response enhancement mechanism for target semantics. By supervising the difference vectors of target semantics between the one-step prediction and the ground truth image, we amplify the effectiveness of target semantics and mitigate the representational gaps inherent in direct cross-modal supervision. Extensive experiments demonstrate that SPF-Portrait achieves state-of-the-art (SOTA) performance in preventing semantic pollution for text-driven portrait customization compared to previous methods.
In summary, our contributions are as follows:
@@
  • We propose SPF-Portrait, a dual-path contrastive learning pipeline, which is the pioneering work to address semantic pollution in text-driven portrait customization.
@@
  • We introduce a novel Semantic-Aware Fine Control alignment process capable of preserving original model’s behavior while meticulously preventing over-alignment.
@@
  • We design a response enhancement mechanism to improve the effectiveness of target semantics while alleviating representation gaps in direct cross-modal supervision.
@@ 2 Related Work
Fine-tuning for T2I Diffusion Models. Numerous solutions [54, 49, 25, 16, 55, 50, 44, 3, 20, 30] have improved the existing T2I diffusion models in various aspects based primarily on fine-tuning [43, 45, 32]. Building upon the fine-tuning paradigm, PEFT-based methods rapidly adapt to new concepts by introducing additional parameters to the original model architecture. LoRA [19] achieves it through low-rank linear layers, while FouRA [3] based on LoRA further improves multi-concept integration by leveraging frequency domain learning. Subsequent studies [14, 40, 34] has further improved the preservation of prior knowledge during fine-tuning. For instance, SVDiff [14] fine-tunes only the singular values, the key parameters, via singular value decomposition. OFT [40] maintains the orthogonality of weight matrices, thereby preserving the hyperspherical energy of the pre-trained model. Although these works preserve pre-trained knowledge while adapting to new concepts, they overlook impure learning from relying solely on diffusion loss, causing new concepts to couple with irrelevant dataset attributes.
@@
Decoupling Generation of Diffusion Models. Efforts have also been made on decoupling control mechanisms, both between image-to-text conditions and within textual conditions, aiming to preventing the hinder to the textual control [58, 7, 39, 20, 13, 5]. To achieve the coupling within text, Magnet [58] and TEBopt [7] analyze and optimize the condition embedding without additional training. However, while mitigating coupling at the instance level, they fail to correct the model’s deviation in understanding refined attributes. RealCustom [20] dynamically adjusts image feature injection based on their impact on diffusion process, while DEADiff [39] tackles similar issues via a decoupling representation mechanism. PuLID [13] employs contrastive learning to prevent the injection of ID from disrupting the textual guidance to achieve decoupling. However, these methods ignore the disruption from text conditions during fine-tuning with reference images.
@@
Distinction with Text-driven Image Editing Methods. The exceptional capability to adhere to base text enables our method to achieve end-to-end image manipulation directly through T2I model, eliminating the need for additional editing pipelines. While integrating text-driven editing methods [12, 44, 43, 9, 48, 26, 24, 4] into the T2I model pipeline can yield results comparable to ours. For a image generated with T2I model, InstructPix2Pix [4] enables precise image manipulation through textual instructions by leveraging a conditioned diffusion model trained on paired image editing datasets. Similarly, DiffusionCLIP [26] and Asyrp [28], inspired by GAN-based methods [1], utilize a local directional CLIP loss between images and text to manipulate specific attributes. However, the task of our work lies in preventing new textual concepts disrupting T2I models, which fundamentally differs from the goal of I2I editing models that focus on image manipulation.
3 Methodology
In this section, we first describe the preliminaries (Sec. 3.1), and then present the dual-path contrastive learning pipeline of our SPF-Portrait (Sec. 3.2). Subsequently, we introduce in detail the process of using the Semantic-Aware Fine-controllable Map to guide alignment with the original model’s behavior (Sec. 3.3) and present the response enhancement mechanism for target semantics (Sec. 3.4).
Figure 1: Visualization of the Attention Map. The salient regions directly reflect response intensity to the target text ”a hat”. Brighter regions indicate higher attention.
Figure 2: The Dual-Path Contrastive Learning Pipeline of SPF-Portrait. The text in blue is the Base text, while those in red is the Target text. Reference Path takes only Base text as input, while Response Path takes complete text (Base text & Target text) as input.
3.1 Preliminary
@@
Diffusion Models. T2I diffusion models can generate images based on text input through a forward diffusion process and a reverse denoising process [18, 43, 45]. The diffusion process follows the Markov chain to gradually transform an image sample $x_{0}$ into noisy samples $x_{1:T}$ by adding Gaussian noise $\epsilon$ over $T$ steps. The denoising process employs a denoising model $\epsilon_{\theta}$ to predict the added noise using $x_{t}$, $t$, and textual conditions $y$ as inputs, where $\theta$ denotes the learnable parameters and $t\in[0,T]$ is the diffusion process timestep. The optimization process can be described as:
$\mathcal{L}_{diff}=\mathbb{E}_{x_{0},\epsilon\thicksim\mathcal{N}(0,1),t}(\|\epsilon-\epsilon_{\theta}(x_{t},t,E)\|_{2}^{2}),$ (1)
@@
Attention Mechanism. Within the Stable Diffusion [43] which we adopted as T2I diffusion models, textual features are integrated into cross-attention layers following the attention mechanism [47] as:
$\left.\left.\left.\left\{\begin{array}[]{l}\mathrm{Attention}(Q,K,V)=\mathrm{Softmax}(\frac{QK^{T}}{\sqrt{d}})V\\ K=W_{K}E;\ \ V=W_{V}E,\end{array}\right.\right.\right.\right.$ (4)
where $Q$ is derived from the UNet image features, $W_{K}$ and $W_{V}$ represent the learnable linear layers.
3.2 Dual-Path Contrastive Learning Pipeline
Naive fine-tuning can adapt a pre-trained T2I model to generate text-customized portrait attributes but fails to achieve incremental learning. We visualize the attention map [47] of target text after naive fine-tuning in Fig. 1 to analyze the reason. It is evident that the response regions of target semantics are extended to unrelated areas, interfering with other attributes that should remain consistent with the original portrait. This is because fine-tuning based solely on diffusion loss focuses on implicitly modeling the joint distribution of data, rather than capturing the independence of semantic attributes. This leads to the coupling of target semantics with redundant features from fine-tuning dataset (Semantic Pollution) and the erroneous expansion of target response area, ultimately disrupting the original behavior.
@@ 3.2 Dual-Path Contrastive Learning Pipeline
To address these issues, we design a dual-path pipeline that leverages contrastive learning [13, 36] to eliminate semantic pollution. As shown in Fig. 2, it first splits the complete prompts into two parts: the base text for original performance and target text for customized attributes, which are encoded by the text encoder as $E_{base}$ and $E_{tar}$. Specifically, the proposed dual paths including: (i) the reference path is initialized from the parameters of pre-trained model. In contrastive learning, it only takes $E_{base}$ as input and remains frozen, serving as a stable reference on behalf of the original model’s performance; and (ii) the response path is initially fine-tuned with conventional method [18] to gain the ability to respond to customized attributes. During contrastive learning, it takes complete text as input including both $E_{base}$ and $E_{tar}$, with only cross-attention layers trainable. By contrastive learning between dual paths, we specifically design a Semantic-Aware Fine Control alignment process to maintain the original model’s behavior and an response enhancement mechanism for target semantics.
3.3 Semantic-Aware Fine Control Alignment
@@
In this section, we provide a detailed presentation of our novel Semantic-Aware Fine Control alignment process. This process first extracts the attention features $\mathcal{F}_{ref}$ and $\mathcal{F}_{res}$ from the reference path and response path. They represent the response of the UNet features $Q_{ref}$ and $Q_{res}$ to the base textual features $E_{base}$, where $Q_{ref}$ and $Q_{res}$ are features from the corresponding UNet’s cross-attention layer in the contrastive paths. The attention features are represented using a variant of the attention mechanism, i.e., Attention $(K,Q,Q)$. By constraining the similarity between the attention features $\mathcal{F}_{ref}$ and $\mathcal{F}_{res}$ from each cross-attention layer, this process encourages the representation of the base text in the response path to approach the behavior of the original model as:
@@
$\left.\left.\left.\left\{\begin{array}[]{l}\mathcal{F}_{ref}=\mathrm{Softmax}(\frac{K_{ref}(E_{base})\ \ Q_{ref}^{T}}{\sqrt{d}})Q_{ref},\\ \mathcal{F}_{res}=\mathrm{Softmax}(\frac{K^{s}_{res}(E_{base})\ \ Q_{res}^{T}}{\sqrt{d}})Q_{res},\\ \mathcal{L}_{\text{text-consistent}}=\sum_{j=1}^{L}\left\|\mathcal{F}^{j}_{ref}-\mathcal{F}^{j}_{res}\right\|_{2},\end{array}\right.\right.\right.\right.$ (8)
@@
where $K_{ref}(E_{base})$ denotes the key of $E_{base}$ in reference path, $K^{s}_{res}(E_{base})=K_{res}([E_{base},E_{tar}])|_{E_{base}}$ indicates the selected keys of $E_{base}$ in response path. $L$ denotes the attention layer number of the denoising model.
@@
To enhance consistency in fine-grained content, we further constrain the UNet features $Q$ from contrastive paths, which contain comprehensive local detail information and global structure [8, 36]. This is formulated as:
$\mathcal{L}_{\text{fine-grained}}=\sum_{j=1}^{L}\left\|Q^{j}_{ref}-Q^{j}_{res}\right\|_{2}.$ (9)
@@
Figure 3: Analysis of Alignment Process. (a) Vanilla aligning results in the over-alignment with original portrait. (b) For the same customization goal, reference image fine-tuning offers a more distinct target response region than T2I fine-tuning.
@@
Although such contrastive alignment effectively prevents the disruption of the original model in reference image-based customization tasks, similar to PuLID [13], this vanilla alignment of intermediate features in text-driven customization suppresses the response intensity of target attributes, as shown in Fig. 3 (a). This causes the results to overly align with the original portrait. The fundamental distinction lies in the learning objectives. As shown in Fig. 3 (b), since the reference image is inherently decoupled from text and represents a more concrete condition, it allows the model to better capture information and distinguish their response regions, thereby reducing the effect of alignment on the response. In contrast, the semantic boundaries between textual concepts are ambiguous, leading to the inhibition of target attributes during the alignment of other text-conditional attributes.
To address this more challenging issue, we propose a Semantic-Aware Fine Control Map (SFCM) which spatially guides the alignment process to be implemented on the appropriate regions, minimizing its impact on the target response. Specifically, during alignment training, the spatial difference in noise predictions between contrastive paths can serve as prior knowledge for target response, forming a soft map $\mathcal{M}$ as:
$\mathcal{M}(E_{base},E_{tar},x_{t},t)=\\ |\epsilon_{\theta}(x_{t},t,E_{base})-\epsilon_{\theta^{{}^{\prime}}}(x_{t},t,[E_{base},E_{tar}])|,$ (10)
@@
where the $\epsilon_{\theta^{{}^{\prime}}}$ and $\epsilon_{\theta}$ represent the prediction in both response and reference paths, respectively, while $\theta^{{}^{\prime}}$ denoting the learnable parameters. As previously analyzed, Semantic Pollution causes the target response regions to diffuse into areas of other attributes, making the noise difference $\mathcal{M}$ unable to precisely characterize the target response regions. Inspired by the insight that if a phrase in base text exhibits low semantic relevance to target text, the regions highlighted by this phrase should be excluded from the $\mathcal{M}$, we design the Semantic-Aware process to refine the soft map. Concretely, for each phrase feature $E_{base}[i],~{}i=\{1,2,\cdots,P\}$and $P$ is the total number of phrase in base text, we compute its mean of the cross-attention maps across all the UNet layers to localize highlighted regions $\overline{A}_{base}[i]$ as:
$\overline{A}_{base}[i]=\frac{1}{L}\sum_{j=1}^{L}{(A_{base}^{j}[i])},$ (11)
@@
where $A_{base}^{j}[i]$ represents the attention map of the $i$-th phrase embedding $E_{base}[i]$ from the $j$-th layer. Subsequently, to quantify the relevance of exclusion, we leverage the representation capabilities of CLIP to calculate the similarity between $E_{tar}$ and each $E_{base}[i]$. We then weight the $\overline{A}_{base}[i]$ based on the similarity, which used to refine the soft map $\mathcal{M}$, as expressed below:
@@
^$\begin{split}\widehat{\mathcal{M}}=\mathcal{M}-&\sum_{i=1}^{P}\overline{A}_{base}[i]\cdot(1-\gamma(i)),\\ \gamma(i)=&D_{CLIP}(E_{base}[i],E_{tar}),\end{split}$ (12)

### v3

# SPF-Portrait: Towards Pure Text-to-Portrait Customization with Semantic Pollution-Free Fine-Tuning

@@ Abstract.
Fine-tuning a pre-trained Text-to-Image (T2I) model on a tailored portrait dataset is the mainstream method for text-to-portrait customization. However, existing methods often severely impact the original model’s behavior (e.g., changes in ID, layout, etc.) while customizing portrait attributes. To address this issue, we propose SPF-Portrait, a pioneering work to purely understand customized target semantics and minimize disruption to the original model. In our SPF-Portrait, we design a dual-path contrastive learning pipeline, which introduces the original model as a behavioral alignment reference for the conventional fine-tuning path. During the contrastive learning, we propose a novel Semantic-Aware Fine Control Map that indicates the intensity of response regions of the target semantics, to spatially guide the alignment process between the contrastive paths. It adaptively balances the behavioral alignment across different regions and the responsiveness of the target semantics. Furthermore, we propose a novel response enhancement mechanism to reinforce the presentation of target semantics, while mitigating representation discrepancy inherent in direct cross-modal supervision. Through the above strategies, we achieve incremental learning of customized target semantics for pure text-to-portrait customization. Extensive experiments show that SPF-Portrait achieves state-of-the-art performance. Project page: https://spf-portrait.github.io/SPF-Portrait/.
Diffusion Model, Text-to-Image, Portrait Generation
copyright: nonejournal: TOGccs: Computing methodologies Computer vision
Figure 1. Left: The paradigm of text-to-portrait customization. Right: Comparison of text-to-portrait customization across various dimensions. The original portrait, representing the original T2I model’s behavior, is based solely on the Base Text as input. The customized portraits are generated by the fine-tuned model and are based on both the Base Text and Target Text as inputs. During the text-to-portrait customization, our SPF-Portrait is able to achieve customized target semantics while maintaining consistency with the original model’s behavior, compared to naive fine-tuning.
1. Introduction
Fine-tuning pre-trained T2I diffusion models (Rombach et al., 2022; Ramesh et al., 2021; Saharia et al., 2022; Esser et al., [n. d.]) offers an efficient approach for text-to-portrait customization (Huang et al., 2023; Han et al., 2024; He et al., 2024), which adapts the models to generate personalized target attributes. However, as shown in Fig. 1, although conventional naive fine-tuning (Rombach et al., 2022) can achieve target semantics, it has a significant impact on the original model’s behavior, such as altering the portrait’s identity, posture, background, etc. This is because, when the model learns the target semantics, the target semantics become entangled with redundant attributes (Hahm et al., 2024) from the fine-tuning dataset. Consequently, while achieving the customized target semantics, the model not only generates the desired attributes but also inadvertently interferes with other original portrait attributes. We refer to this phenomenon as ”Semantic Pollution”, which is detrimental and often ignored. This further indicates a non-incremental learning. To address this issue, we propose SPF-Portrait, the first method to our knowledge that purely understands customized target semantics while eliminating semantic pollution in text-to-portrait customization. As shown in Fig. 1, our method is capable of stably performing well in customizing portrait attributes across various dimensions.
@@ 1. Introduction
One line of previous research related to mitigating semantic pollution is PEFT-based methods (Hu et al., 2021; Ding et al., 2023; Zhang et al., 2023a; Borse et al., 2024; Liu et al., 2023). They minimize influence through low-rank adapter (e.g., LoRA and its variants (Ding et al., 2023; Zhang et al., 2023a; Borse et al., 2024)) or orthogonal constraints (Qiu et al., 2023; Liu et al., 2023). However, their reliance on diffusion loss for implicit joint distribution modeling (Song et al., 2020), rather than understanding disentangled semantics, only allows for limited preservation of the original behavior. Another line of work (Chefer et al., 2023; Zhuang et al., 2024; Chen et al., 2024; Cai et al., 2024; Mañas et al., 2024; Liu et al., 2024; Jiang et al., 2024) aims to purify the understanding of text embeddings and decouple attributes from each other. They enhance attribute independence through embedding-level decoupling (Magnet (Zhuang et al., 2024), TEBopt (Chen et al., 2024)) or attention regularization (Tokencompose (Wang et al., 2024b)). While effective for instance-level generation (e.g., a cat or a dog), these methods fail when comes to refined attributes, such as hairstyles and skin textures.
@@
Pure text-to-portrait customization manifests itself in generated portraits as introducing differences only by target attributes while maintaining consistency in unrelated attributes with the original model’s outputs. It requires achieving the following two objectives: 1) Effective adaptation of T2I models to target attributes, and 2) Faithful preservation of the original model’s behavior. To this end, we propose the SPF-Portrait that incorporates an additional training stage after naive fine-tuning. In this stage, we design a dual-path contrastive learning pipeline that introduces the frozen original model as the anchor of original behavior for the conventional fine-tuning path. During contrastive learning, we extract and constrain variant attention features and UNet features from the corresponding cross-attention layers in contrastive paths to align with the original performance. We propose a novel Semantic-Aware Fine Control Map (SFCM) that accurately identifies the response regions of target semantics to spatially guide the alignment of these intermediate features. This alignment process precisely aligns irrelevant attributes, avoiding suppression of target attribute and over-alignment. Moreover, we propose a response enhancement mechanism for target semantics. By supervising the difference vectors of target semantics between the one-step prediction and the ground truth image, we mitigate the representational gaps inherent in direct cross-modal supervision and enhance the manifestation of target semantics. Extensive experiments show that SPF-Portrait achieves state-of-the-art performance in preventing semantic pollution for pure text-to-portrait customization. In summary, our contributions are as follows:
@@
  • We propose SPF-Portrait, a dual-path contrastive learning pipeline, which is the pioneering work to address semantic pollution in text-to-portrait customization.
@@
  • We introduce a novel Semantic-Aware Fine Control alignment process capable of preserving the original model’s behavior while meticulously preventing over-alignment.
@@
  • We design a response enhancement mechanism to improve the presentation of target semantics while alleviating representation gaps in direct cross-modal supervision.
@@ 2. Related Work
Fine-tuning for T2I Diffusion Models. Numerous solutions (Zhang et al., 2023b; Wang et al., 2024a; Ruiz et al., 2023; Huang et al., 2024; Li et al., 2024b; Liao et al., 2024) have improved existing T2I diffusion models (Rombach et al., 2022; Lin et al., 2024) in various aspects based primarily on fine-tuning. Building upon the fine-tuning paradigm (Liao et al., 2025; Luo et al., 2024, 2025; Wan et al., 2024), PEFT-based methods (Wu et al., 2024; Zhang et al., 2023a; Borse et al., 2024) rapidly adapt to new concepts by introducing additional parameters to the original model. LoRA (Hu et al., 2021) achieves this through low-rank linear layers, while FouRA (Borse et al., 2024) based on LoRA further improves multi-concept integration by leveraging frequency domain learning. Subsequent studies (Han et al., 2023; Qiu et al., 2023; Liu et al., 2023) further improve the preservation of prior knowledge during fine-tuning. For instance, SVDiff (Han et al., 2023) fine-tunes only the singular values, the key parameters, via singular value decomposition. OFT (Qiu et al., 2023) maintains the orthogonality of weight matrices, thereby preserving the hyperspherical energy of the pre-trained model. Although they preserve pre-trained knowledge while adapting to new concepts, they overlook impure learning from relying solely on diffusion loss, causing new attributes to couple with irrelevant dataset attributes.
@@
Decoupling Generation of Diffusion Models. Efforts have also been made on decoupling control mechanisms, both between image-to-text conditions and within textual conditions, aiming to preventing the hinder to the textual control (Zhuang et al., 2024; Chen et al., 2024; Qi et al., 2024; Huang et al., 2024; Xing et al., 2024; Gao et al., 2024; Chang et al., 2024). To achieve the coupling within text, Magnet (Zhuang et al., 2024) and TEBopt (Chen et al., 2024) analyze and optimize the condition embedding without additional training. However, while mitigating coupling at the instance level, they fail to correct the model’s deviation in understanding refined attributes. RealCustom (Huang et al., 2024) dynamically adjusts image feature injection based on their impact on diffusion process, while DEADiff (Qi et al., 2024) tackles similar issues via a decoupling representation mechanism. PuLID (Guo et al., 2024) employs contrastive learning to prevent the injection of ID from disrupting the textual guidance to achieve decoupling. However, these methods ignore the disruption from text conditions during fine-tuning with reference images.
@@
Distinction with Text-driven Image Editing Methods. The exceptional capability to adhere to base text enables our method to achieve end-to-end image manipulation (Brack et al., 2024; Hoogeboom et al., 2023; Gan et al., 2023) directly through T2I model, eliminating the need for additional editing pipelines. While integrating text-driven editing methods (Deutch et al., 2024; Wang et al., 2024c; Kim et al., 2022; Ju et al., 2024; Brooks et al., 2023) into the T2I model pipeline can yield results comparable to ours. For a image generated with T2I model, InstructPix2Pix (Brooks et al., 2023) enables precise image manipulation through textual instructions by leveraging a conditioned diffusion model trained on paired image editing datasets. Similarly, DiffusionCLIP (Kim et al., 2022) and Asyrp (Kwon et al., 2022), inspired by GAN-based methods (Alaluf et al., 2022), utilize a local directional CLIP loss (Baykal et al., 2023) between images and text to manipulate specific attributes. However, the task of our work lies in preventing new textual attributes from disrupting T2I models, which fundamentally differs from the goal of I2I editing models that focus on image manipulation.
3. Methodology
Our SPF-Portrait improves naive fine-tuning by introducing an additional training stage. In the first stage, we employ naive fine-tuning to strive for the preliminary response to target semantics without considering the contamination to the original model. In the second stage, we design a Dual-path Contrastive Learning approach (Sec. 3.2) that introduces the frozen original model along with the fine-tuning path. During contrastive learning, we propose the Semantic-Aware Fine-Control Map to guide alignment with the original model’s behavior (Sec. 3.3) and design the Response Enhancement mechanism for target semantics (Sec. 3.4).
3.1. Preliminary
@@
Diffusion Models. T2I diffusion models generate images based on text input through a forward diffusion process and a reverse denoising process (Ho et al., 2020; Saharia et al., 2022). The diffusion process follows the Markov chain to transform an image sample $x_{0}$ into noisy samples $x_{1:T}$ by adding Gaussian noise $\epsilon$ over $T$ steps. The denoising process employs a denoising model $\epsilon_{\theta}$ to predict the added noise using $x_{t}$, $t$, and textual conditions $y$ as inputs, where $\theta$ denotes the learnable parameters and $t\in[0,T]$ is the diffusion process timestep. The optimization process can be described as:
(1) $\mathcal{L}_{diff}=\mathbb{E}_{x_{0},\epsilon\thicksim\mathcal{N}(0,1),t}(\|\epsilon-\epsilon_{\theta}(x_{t},t,E)\|_{2}^{2}),$
@@
3.2. Dual-Path Contrastive Learning Pipeline
@@
Although the first stage of training, a naive fine-tuning, can initially achieve the adaptation of T2I models to target attributes. However, as shown in Fig. 1, it will severely affect the behavior of the original model. We visualize the attention map (Vaswani, 2017) of target text after naive fine-tuning in Fig. 2 to diagnose this limitation. The response regions of the target semantics are extended to unrelated areas, interfering with other attributes, which is caused by semantic pollution during fine-tuning. To address this issue, we design an additional training stage that utilizes a dual-path contrastive learning pipeline. Specifically, the proposed dual paths including: (i) Reference Path comprises a frozen model initialized from the original pre-trained T2I model. In contrastive learning, it only takes $E^{ref}_{base}=\tau_{text}(y_{base})$ as input, serving as a stable reference on behalf of the original model’s behavior; and (ii) Response Path includes a model initially resumed from the first stage. During contrastive learning stage, it takes complete text (i.e., $y_{base}$ and $y_{tar}$) as input:
(2) $\begin{split}E_{base}^{res}&=\tau_{text}([y_{base},y_{tar}])|_{y_{base}},\\ E_{tar}&=\tau_{text}([y_{base},y_{tar}])|_{y_{tar}},\end{split}$
where $[y_{base},y_{tar}]$ represents the concatenated text prompt. $E_{base}^{res}$ and $E_{tar}$ represent the encoded feature segments corresponding to $y_{base}$ and $y_{tar}$ portions respectively. By contrastive learning between dual paths, we specifically design a Semantic-Aware Fine Control alignment process to maintain the original model’s behavior and an response enhancement mechanism for target semantics.
Figure 2. Visualization of the Attention Map. The salient regions directly reflect response intensity to the target semantics ”a hat”.
Figure 3. The Dual-Path Contrastive Learning Pipeline of SPF-Portrait. The text in blue is the Base text, while those in red is the Target text. Reference Path takes only Base text as input, while Response Path takes complete text (Base text & Target text) as input.
3.3. Semantic-Aware Fine Control Alignment
@@
In this section, we provide a detailed presentation of our novel Semantic-Aware Fine Control alignment process. This process first extracts the attention features $\mathcal{F}_{ref}$ and $\mathcal{F}_{res}$ from the reference path and response path. These features are derived from a variant of the standard attention mechanism, i.e., Attention $(K,Q,Q)$. They represent the response of the UNet features $Q_{ref}$ and $Q_{res}$ to the base textual features $E_{base}$, where $Q_{ref}$ and $Q_{res}$ are features from the corresponding UNet’s cross-attention layer in the contrastive paths. By constraining the similarity between the attention features $\mathcal{F}_{ref}$ and $\mathcal{F}_{res}$ from each cross-attention layer, this process encourages the representation of the base text in the response path to approach the behavior of the original model as:
@@
(6) $\left.\left.\left.\left\{\begin{array}[]{l}\mathcal{F}_{ref}=\mathrm{Softmax}(\frac{K_{ref}(E^{ref}_{base})\ \ Q_{ref}^{T}}{\sqrt{d}})Q_{ref},\\ \mathcal{F}_{res}=\mathrm{Softmax}(\frac{K_{res}(E^{res}_{base})\ \ Q_{res}^{T}}{\sqrt{d}})Q_{res},\\ \mathcal{L}_{\text{text-consistent}}=\sum_{j=1}^{L}\left\|\mathcal{F}^{j}_{ref}-\mathcal{F}^{j}_{res}\right\|_{2},\end{array}\right.\right.\right.\right.$
@@
where $K_{ref}$ and $K_{res}$ denotes the key of $E^{ref}_{base}$ and $E^{res}_{base}$ in dual-path. $L$ represents the attention layer number of the denoising model.
@@
To enhance consistency in fine-grained content, we further constrain the UNet features $Q$ from contrastive paths, which contains comprehensive information on local details and global structure (Chung et al., 2024; Mo et al., 2024). This is formulated as:
(7) $\mathcal{L}_{\text{fine-grained}}=\sum_{j=1}^{L}\left\|Q^{j}_{ref}-Q^{j}_{res}\right\|_{2}.$
@@
Figure 4. Analysis of Alignment Process. (a) Vanilla alignment results in the over-alignment with original portrait. (b) For the same customization attribute, reference image-based fine-tuning offers a more distinct target response region than T2I fine-tuning.
@@
Although such a contrastive alignment effectively prevents the impact of the original model (e.g., in reference image-based customization tasks (Guo et al., 2024)), this vanilla alignment of intermediate features in text-driven generation suppresses the response intensity of target semantics, as shown in Fig. 4 (a). This causes the customized portrait to overly align with the original portrait. As shown in Fig. 4 (b), the fundamental distinction lies in the learning objectives. Since the reference image is inherently decoupled from text and represents a more concrete condition, it allows the model to have well-defined objectives to consult, thereby having negligible disturbance on target attribute performance. In contrast, the semantic boundaries between textual concepts are ambiguous, which finally amplifies the influence. To address this more challenging issue, we propose a Semantic-Aware Fine Control Map (SFCM) that spatially guides the alignment process to be implemented in the appropriate regions, minimizing its disturbance on the target response. Specifically, during alignment training, the spatial difference in noise predictions between contrastive paths can serve as prior knowledge for target response, forming a soft map $\mathcal{M}$ as:
(8) $\mathcal{M}=|\epsilon_{\theta}(x_{t},t,E^{ref}_{base})-\epsilon_{\theta^{{}^{\prime}}}(x_{t},t,[E^{res}_{base},E_{tar}])|,$
@@
where the $\epsilon_{\theta^{{}^{\prime}}}$ and $\epsilon_{\theta}$ represent the prediction in both response and reference paths, respectively, while $\theta^{{}^{\prime}}$ denoting the learnable parameters. As previously analyzed, Semantic Pollution causes the target response regions to diffuse into areas of other attributes, making the noise difference $\mathcal{M}$ unable to precisely characterize the target response regions. Inspired by the insight that if a phrase in base text exhibits low semantic relevance to target text, the regions highlighted by this phrase should be excluded from the $\mathcal{M}$, we design the Semantic-Aware process to refine the soft map. For the input base text in response path, we split it into multiple phrases, as shown in ”Phrase embedding Attention Map” of Fig. 3. Concretely, for each phrase feature $E^{res}_{base}[i],~{}i=\{1,2,\cdots,P\}$and $P$ is the total number of phrase in base text, we compute its mean of the cross-attention maps across all the UNet layers to localize highlighted regions $\overline{A}_{base}[i]$ as:
(9) $\overline{A}_{base}[i]=\frac{1}{L}\sum_{j=1}^{L}{(A_{base}^{j}[i])},$
@@
where $A_{base}^{j}[i]$ represents the attention map of the $i$-th phrase embedding $E^{res}_{base}[i]$ from the $j$-th layer. Subsequently, to quantify the relevance of exclusion, we leverage the representation capabilities of CLIP to calculate the similarity between $E_{tar}$ and each $E^{res}_{base}[i]$. We then weight the $\overline{A}_{base}[i]$ based on the similarity, which used to refine the soft map $\mathcal{M}$, as expressed below:
@@
(10) ^$\begin{split}\widehat{\mathcal{M}}=\mathcal{M}-&\sum_{i=1}^{P}\overline{A}_{base}[i]\cdot(1-\gamma(i)),\\ \gamma(i)=&D_{CLIP}(E^{res}_{base}[i],E_{tar}),\end{split}$